an electron enterd uniform field that is directed downward and has a magnitude of 5 N/C. (a.)find the magnitude and direction of the force experienced by the electron (b.) find also its acceleration


1
Expert's answer
2022-02-28T10:40:18-0500

(a) find the magnitude and direction of the force experienced by the electron

"F=eE=1.6*10^{-19}*5=8*10^{-19}\:\rm N"

The force is directed upward.

(b) find also its acceleration

"a=\frac{F}{m}=\frac{8*10^{-19}}{9.1*10^{-31}}=8.8*10^{11}\:\rm m/s^2"